OpenStudy (anonymous):

a circle with the center at the origin and a radius of 8 units

OpenStudy (anonymous):

The equation of the circle is: x^2 + y^2 = 64

OpenStudy (anonymous):

\[x^2+y^2-14x-2y+25=0\]

OpenStudy (anonymous):

\[x^2+y^2-14x-2y+25=0\]